# Position Paper: Against Spurious Sparks—Dovelating Inflated AI Claims [Uncaptioned image]

### v2

@@ Abstract
Humans have a tendency to see ‘human’-like qualities in objects around them. We name our cars, and talk to pets and even household appliances, as if they could understand us as other humans do. This behavior, called anthropomorphism, is also seeing traction in Machine Learning (ML), where human-like intelligence is claimed to be perceived in Large Language Models (LLMs). In this position paper, considering professional incentives, human biases, and general methodological setups, we discuss how the current search for Artificial General Intelligence (AGI) is a perfect storm for over-attributing human-like qualities to LLMs. In several experiments, we demonstrate that the discovery of human-interpretable patterns in latent spaces should not be a surprising outcome. Also in consideration of common AI portrayal in the media, we call for the academic community to exercise extra caution, and to be extra aware of principles of academic integrity, in interpreting and communicating about AI research outcomes.
**footnotetext: These authors contributed equally to this work.
Keywords Machine Learning, Anthropomorphism, Artificial General Intelligence
@@ 1 Introduction
At the same time, in the way in which academic research is disseminated in the AI and machine learning fields today, major shifts are happening. Where these research fields have actively adopted early sharing of preprints and code, the volume of publishable work has exploded to a degree that one cannot reasonably keep up with broad state-of-the-art, and social media influencers start playing a role in article discovery and citeability (Weissburg et al., 2024). Furthermore, because of major commercial stakes with regard to AI and machine learning technology, and e.g. following the enthusiastic societal uptake of products employing large language models (LLMs), such as ChatGPT, the pressure to beat competitors as fast as possible is only increasing, and strong eagerness can be observed in many domains to ‘do something with AI’ in order to innovate and remain current.
@@
Generally, in the popularity of ChatGPT and the integration of generative AI in productivity tools (e.g. through Microsoft’s Copilot integrations in GitHub and Office applications), one also can wonder whether the promise of AI is more in computationally achieving general intelligence, or rather in the engineering of general-purpose tools [^1] A Swiss army knife is an effective general-purpose tool, without people wondering whether it is intelligent.. Regardless, stakes and interests are high, e.g. with ChatGPT clearing nearly \$1 billion in months of its release [^2] https://www.bloomberg.com/news/articles/2023-08-30/openai-nears-1-billion-of-annual-sales-as-chatgpt-takes-off.
@@
When combining massive financial incentives with the presence of a challenging and difficult-to-understand technology, that aims towards human-like problem-solving and communication abilities, a situation arises that is fertile for the misinterpretation of spurious cues as hints towards AGI, or other qualities like sentience [^3] https://www.scientificamerican.com/article/google-engineer-claims-ai-chatbot-is-sentient-why-that-matters/ and consciousness. AI technology only becomes more difficult to understand as academic publishing in the space largely favors performance, generalization, quantitative evidence, efficiency, building on past work, and novelty (Birhane et al., 2022). As such, works that make it into top-tier venues tend to propose heavier and more complicated technical takes on tasks that (in the push towards generalizability) get more vague, while the scaling-up of data makes traceability of possible memorization harder. In a submission-overloaded reality, researchers may further get incentivized to oversell and overstate achievement claims.
@@
Noticing these trends, we as the authors of this article are concerned. We feel that the current culture of racing toward Big Outcome Statements in industry and academic publishing too much disincentivizes efforts toward more thorough and nuanced actual problem understanding. At the same time, as the outside world is so eager to adopt AI technology, (too) strong claims make for good sales pitches, but a question is whether there is indeed sufficient evidence for these claims. With successful A(G)I outcomes needing to look human-like, this also directly plays into risks of anthropomorphizing (the attribution of human-like qualities to non-human objects) and confirmation bias (the seeking-out and/or biased interpretation of evidence in support of one’s beliefs). In other words, it is very tempting to claim surprising human-like achievements of AI, and as humans, we are very prone to genuinely believing this.
@@
To strengthen our argument, in this paper, we first consider a recently viral work (Gurnee and Tegmark, 2023a) in which claims about the learning of world models by LLMs were made. We present several experiments that may invite similar claims on models yielding more intelligent outcomes than would have been expected—while at the same time indicating how we feel these claims should not be made. Furthermore, we present a review of social science findings that underline how prone humans are to being enticed by patterns that are not really there. Combining this with the way in which media portrayal of AI has tended towards science-fiction imagery of mankind-threatening robots, we argue that the current AI culture is a perfect storm for making and believing inflated claims, and call upon our fellow academics to be extra mindful and scrutinous about this.
@@ 2 Surprising Patterns in Latent Spaces?
In 2023, a research article went viral on the X [^4] https://twitter.com/wesg52/status/1709551516577902782?s=20 platform (Gurnee and Tegmark, 2023b). Through linear probing experiments, the claim was made that LLMs learned literal maps of the world. As such, they were considered to be more than ‘stochastic parrots’ (Bender et al., 2021) that can only correlate and mimic existing patterns from data, but not truly understand it. While the manuscript immediately received public criticism (Marcus, 2023), and the revised, current version is more careful with regard to its claims (Gurnee and Tegmark, 2023a), reactions on X seemed to largely exhibit excitement and surprise at the authors’ findings. However, in this section, through various simple examples, we make the point that observing patterns in latent spaces should not be a surprising revelation. After starting with a playful example of how easy it is to ‘observe’ a world model, we build up a larger example focusing on key economic indicators and central bank communications.
@@ 2.1 Are Neural Networks Born with World Models?
Gurnee and Tegmark (2023a) extract and visualize the alleged geographical world model by training linear regression probes on internal activations in LLMs (including Llama-2) for the names of places, to predict geographical coordinates associated with these places. Now, the Llama-2 model has ingested huge amounts of publicly available data from the internet, including Wikipedia dumps from the June-August 2022 period (Touvron et al., 2023). It is therefore highly likely that the training data contains geographical coordinates, either directly or indirectly. At the very least, we should expect that the model has seen features during training that are highly correlated with geographical coordinates. The model itself is essentially a very large latent space to which all features are randomly projected in the very first instance before being passed through a series of layers which are gradually trained for downstream tasks.
In our first example, we simulate this scenario, stopping short of training the model. In particular, we take the world_place.csv that was used in Gurnee and Tegmark (2023a), which maps locations/areas to their latitude and longitude. For each place, it also indicates the corresponding country. From this, we take the subset that contains countries that are currently part of the top 10 FIFA world ranking, and assign the current rank to each country (i.e., Argentina gets 1, France gets 2, …). To ensure that the training data only involves a noisy version of the coordinates, we transform the longitude and latitude data as follows: $\rho\cdot\text{coord}+(1-\rho)\cdot\epsilon$ where $\rho=0.5$ and $\epsilon\sim\mathcal{N}(0,5)$.
@@
While the fit certainly is not perfect, the results do indicate that the random projection contains representations that are useful for the task at hand. Thus, this simple example illustrates that meaningful target representations should be recoverable from a sufficiently large latent space, given the projection of a small number of highly correlated features. Similarly, Alain and Bengio (2016) observe that even before training a convolutional neural network on MNIST data, the layer-wise activations can already be used to perform binary classification. In fact, it is well-known that random projections can be used for prediction tasks (Dasgupta, 2013).
@@ 2.2 PCA as a Yield Curve Interpreter
We now move to a concrete application domain: Economics. Here, the yield curve, plotting the yields of bonds against their maturities, is a popular tool for investors and economists to gauge the health of the economy. The yield curve’s slope is often used as a predictor of future economic activity: a steep yield curve is associated with a growing economy, while a flat or inverted yield curve is associated with a contracting economy. To leverage this information in downstream modelling tasks, economists regularly use PCA to extract a low-dimensional projection of the yield curve that captures relevant variation in the data (e.g. Berardi and Plazzi (2022), Kumar (2022) and Crump and Gospodinov ).
@@ 2.3 LLMs for Economic Sentiment Prediction
Closely following the approach in Gurnee and Tegmark (2023a), we apply it to the novel Trillion Dollar Words (Shah et al., 2023) financial dataset, containing a curated selection of sentences formulated by central bankers of the US Federal Reserve, and communicated to the public in speeches, meeting minutes and press conferences.  (Shah et al., 2023) use this dataset to train a set of LLMs and rule-based models to classify sentences as either ‘dovish’, ‘hawkish’ or ‘neutral’. In the context of central banking, ‘hawkishness’ is typically associated with tight monetary policy: in other words, a ‘hawkish’ stance on policy favors high interest rates to limit the supply of money and thereby control inflation. The authors first manually annotate a sub-sample of the available data and then fine-tune various models for the classification task. Their model of choice, FOMC-RoBERTa (a fine-tuned version of RoBERTa (Liu et al., 2019)), achieves an $F_{1}$ score of around $>0.7$ on the test data. To illustrate the potential usefulness of the learned classifier, they use predicted labels for the entire dataset to compute an ad-hoc, count-based measure of ‘hawkishness’. This measure is shown to correlate with key economic indicators in the expected direction: when inflationary pressures rise, the measured level of ‘hawkishness’ increases, as central bankers react by raising interest rates to bring inflation back to target.
@@ 2.3.1 Linear Probes
For each layer, we compute linear probes through Ridge regression on two inflation indicators (the Consumer Price Index (CPI) and the Producer Price Index (PPI)) and US Treasury yields at different levels of maturity. To allow comparison with Shah et al. (2023), we let yields enter the regressions in levels. To measure price inflation we use percentage changes proxied by log differences. To mitigate issues related to over-parameterization, we follow the recommendation in Alain and Bengio (2016) to first reduce the dimensionality of the computed activations.In particular, we restrict our linear probes to the first 128 principal components of the embeddings of each layer. To account for stochasticity, we use an expanding window scheme with 5 folds for each indicator and layer. To avoid look-ahead bias, PCA is always computed on the sub-samples used for training the probe.
Figure 3 shows the out-of-sample root mean squared error (RMSE) for the linear probe, plotted against FOMC-RoBERTa’s $n$-th layer. The values correspond to averages across cross-validation folds. Consistent with related work (Alain and Bengio, 2016; Gurnee and Tegmark, 2023a), we observe that model performance tends to be higher for layers near the end of the transformer model. Curiously, for yields at longer maturities, we find that performance eventually deteriorates for the very final layers. We do not observe this for the training data, so we attribute this to overfitting.
@@
Figure 3: Out-of-sample root mean squared error (RMSE) for the linear probe plotted against FOMC-RoBERTa’s $n$-th layer for different indicators. The values correspond to averages computed across cross-validation folds, where we have used an expanding window approach to split the time series.
@@ 2.3.2 Sparks of economic understanding?
Even though FOMC-RoBERTa, which is substantially smaller than the models tested in Gurnee and Tegmark (2023a), was not explicitly trained to uncover associations between central bank communications and the level of consumer prices, it appears that the model has distilled representations that can be used to predict inflation (although they certainly will not win any forecasting competitions). So, have we uncovered further evidence that LLMs “aren’t mere stochastic parrots”? Has FOMC-RoBERTa developed an intrinsic ‘understanding’ of the economy just by ‘reading’ central bank communications? Thus, can economists readily adopt FOMC-RoBERTa as a domain-relevant tool?
@@ 3 Human Proneness to Over-Interpretation
Linear probes and related tools from mechanistic interpretability were proposed in the context of monitoring models and diagnosing potential problems (Alain and Bengio, 2016). Favorable outcomes from probes merely indicate that the model “has learned information relevant for the property [of interest]" (Belinkov, 2021). Our examples demonstrate that this is achievable even for small models, while these have certainly not developed intrinsic “understanding” of the world. Thus, we argue that more conservative and rigorous tests for emerging capabilities of AI model are needed.
Generally, humans are prone to seek patterns everywhere. Meaningful patterns have proven useful in helping us make sense of our past, navigate our present and predict the future. Although this tendency to perceive patterns likely leads to evolutionary benefits even when the perceived patterns are false (Foster and Kokko, 2009), psychology has revealed a host of situations in which the ability to perceive patterns severely misfires, leading to irrational beliefs in e.g. the power of superstitions (Foster and Kokko, 2009), conspiracy theories (Van Prooijen et al., 2018), the paranormal (Müller and Hartmann, 2023), gambler’s fallacies (Ladouceur et al., 1996) and even interpreting pseudo-profound bullshit as meaningful (Walker et al., 2019).

### v3

# Position: Stop Making Unscientific AGI Performance Claims

@@ Abstract
Developments in the field of Artificial Intelligence (AI), and particularly large language models (LLMs), have created a ‘perfect storm’ for observing ‘sparks’ of Artificial General Intelligence (AGI) that are spurious. Like simpler models, LLMs distill meaningful representations in their latent embeddings that have been shown to correlate with external variables. Nonetheless, the correlation of such representations has often been linked to human-like intelligence in the latter but not the former. We probe models of varying complexity including random projections, matrix decompositions, deep autoencoders and transformers: all of them successfully distill information that can be used to predict latent or external variables and yet none of them have previously been linked to AGI. We argue and empirically demonstrate that the finding of meaningful patterns in latent spaces of models cannot be seen as evidence in favor of AGI. Additionally, we review literature from the social sciences that shows that humans are prone to seek such patterns and anthropomorphize. We conclude that both the methodological setup and common public image of AI are ideal for the misinterpretation that correlations between model representations and some variables of interest are ‘caused’ by the model’s understanding of underlying ‘ground truth’ relationships. We, therefore, call for the academic community to exercise extra caution, and to be keenly aware of principles of academic integrity, in interpreting and communicating about AI research outcomes.
Machine Learning, Anthropomorphism, Artificial General Intelligence
@@ 1 Introduction
At the same time, in the way in which academic research is disseminated in the AI and machine learning fields today, major shifts are happening. Where these research fields have actively adopted early sharing of preprints and code, the volume of publishable work has exploded to a degree that one cannot reasonably keep up with broad state-of-the-art, and social media influencers start playing a role in article discovery and citeability (Weissburg et al., 2024). Furthermore, because of major commercial stakes with regard to AI and machine learning technology, and e.g. following the enthusiastic societal uptake of products employing LLMs, such as ChatGPT, the pressure to beat competitors as fast as possible is only increasing, and strong eagerness can be observed in many domains to ‘do something with AI’ in order to innovate and remain current.
@@
Generally, in the popularity of ChatGPT and the integration of generative AI in productivity tools (e.g. through Microsoft’s Copilot integrations in GitHub and Office applications), one also can wonder whether the promise of AI is more in computationally achieving general intelligence, or rather in the engineering of general-purpose tools [^1] A Swiss army knife is an effective general-purpose tool, without people wondering whether it exhibits intelligence.. Regardless, stakes and interests are high, e.g. with ChatGPT clearing nearly \$1 billion in months of its release [^2] https://www.bloomberg.com/news/articles/2023-08-30/openai-nears-1-billion-of-annual-sales-as-chatgpt-takes-off.
@@
When combining massive financial incentives with the presence of a challenging and difficult-to-understand technology, that aims towards human-like problem-solving and communication abilities, a situation arises that is fertile for the misinterpretation of spurious cues as hints towards AGI, or other qualities like sentience [^3] https://www.scientificamerican.com/article/google-engineer-claims-ai-chatbot-is-sentient-why-that-matters/ and consciousness. AI technology only becomes more difficult to understand as academic publishing in the space largely favors performance, generalization, quantitative evidence, efficiency, building on past work, and novelty (Birhane et al., 2022). As such, works that make it into top-tier venues tend to propose heavier and more complicated technical takes on tasks that (in the push towards generalizability) get more vague, while the scaling-up of data makes traceability of possible memorization harder. In a submission-overloaded reality, researchers may further get incentivized to oversell and overstate achievement claims. At the same time, while currently popular in literature, inherent complexity and opaqueness in technical solutions may fundamentally be unwise to pursue in high-stakes applications (Rudin, 2019).
@@
Noticing these trends, we as the authors of this article are concerned. We feel that the current culture of racing toward Big Outcome Statements in industry and academic publishing too much disincentivizes efforts toward more thorough and nuanced actual problem understanding. At the same time, as the outside world is so eager to adopt AI technology, (too) strong claims make for good sales pitches, but a question is whether there is indeed sufficient evidence for these claims. With successful AGI outcomes needing to look human-like, this also directly plays into risks of anthropomorphizing (the attribution of human-like qualities to non-human objects) and confirmation bias (the seeking-out and/or biased interpretation of evidence in support of one’s beliefs). In other words, it is very tempting to claim surprising human-like achievements of AI, and as humans, we are very prone to genuinely believing this. We therefore urge our fellow researchers to stop making unscientific AGI performance claims.
@@
To strengthen our argument, in this paper, we first present related work in Section 2. We then consider a recently viral work (Gurnee & Tegmark, 2023b) in which claims about the learning of world models by LLMs were made. In Section 3, we present several experiments that may invite similar claims on models yielding more intelligent outcomes than would have been expected—while at the same time indicating how we feel these claims should not be made. Furthermore, we present a review of social science findings in Section 4 that underline how prone humans are to being enticed by patterns that are not really there. Combining this with the way in which media portrayal of AI has tended towards science-fiction imagery of mankind-threatening robots, we argue that the current AI culture is a perfect storm for making and believing inflated claims, and call upon our fellow academics to be extra mindful and scrutinous about this. Finally, in Section 5, we propose specific structural and cultural changes to improve the current situation. Section 6 concludes.
2 Related Work
In this work, we question the practice of using outcomes from mechanistic interpretability to support AGI claims. This is not to be seen as criticism toward the underlying methodologies in isolation, but rather in the context of current publishing practices that we repeatedly challenge throughout this work. Many closely related works are free of any grandiose conclusions and instead highlight the benefits of mechanistic interpretability that we also highlight here (Nanda et al., 2023; Gurnee et al., 2023; Li et al., 2022).
Another related subfield investigates the capacity of LLMs to reason causally. Here, too, there is an opportunity to over-interpret the finding of causal information as causal understanding. Recent work has shown, for example, that LLMs can indeed correctly predict causal relationships and this may have practical use cases (Kıcıman et al., 2023). But despite the potential utility, the authors also demonstrate that this capacity can be partially explained by memorization, rather than an actual understanding of causal relationships. Similarly, Zečević et al. (2023) provide evidence indicating that current LLMs “may talk causality but are not causal”.
Two other recent works are related to this work and align well with the position we present here. Schaeffer et al. (2024) demonstrate that the apparent emergent abilities of large language models may be driven by a choice of evaluation metrics, rather than some fundamental property that is intrinsic to this family of models. Their work highlights the need for rigorous testing and benchmarking of LLMs, which we also point to in this work, albeit in a slightly different methodological context. Kloft et al. (2024) provide experimental evidence demonstrating that people have heightened expectations and a biased, positive view of AI. The authors run a user study of human-AI interaction, in which participants performed better at a given task when they (wrongly) thought they were aided by a positively described AI. This placebo effect was found to be robust to negative descriptions of AI. Positive bias towards AI may exacerbate other factors that drive people to make unscientific claims about the current state of AI, which we discuss in Section 4.
@@ 3 Surprising Patterns in Latent Spaces?
In 2023, a research article went viral on the X [^4] https://twitter.com/wesg52/status/1709551516577902782?s=20 platform (Gurnee & Tegmark, 2023a). Through linear probing experiments, the claim was made that LLMs learned literal maps of the world. As such, they were considered to be more than ‘stochastic parrots’ (Bender et al., 2021) that can only correlate and mimic existing patterns from data, but not truly understand it. While the manuscript immediately received public criticism (Marcus, 2023), and the revised, current version is more careful with regard to its claims (Gurnee & Tegmark, 2023b), reactions on X seemed to largely exhibit excitement and surprise at the authors’ findings. However, in this section, through various simple examples, we make the point that observing patterns in latent spaces should not be a surprising revelation. After starting with a playful example of how easy it is to ‘observe’ a world model, we build up a larger example focusing on key economic indicators and central bank communications.
@@ 3.1 Are Neural Networks Born with World Models?
Gurnee & Tegmark (2023b) extract and visualize the alleged geographical world model by training linear regression probes on internal activations in LLMs (including Llama-2) for the names of places, to predict geographical coordinates associated with these places. Now, the Llama-2 model has ingested huge amounts of publicly available data from the internet, including Wikipedia dumps from the June-August 2022 period (Touvron et al., 2023). It is therefore highly likely that the training data contains geographical coordinates, either directly or indirectly. At the very least, we should expect that the model has seen features during training that are highly correlated with geographical coordinates. The model itself is essentially a very large latent space to which all features are randomly projected in the very first instance before being passed through a series of layers which are gradually trained for downstream tasks.
In our first example, we simulate this scenario, stopping short of training the model. In particular, we take the world_place.csv that was used in Gurnee & Tegmark (2023b), which maps locations/areas to their latitude and longitude. For each place, it also indicates the corresponding country. From this, we take the subset that contains countries that are currently part of the top 10 FIFA world ranking, and assign the current rank to each country (i.e., Argentina gets 1, France gets 2, …). To ensure that the training data only involves a noisy version of the coordinates, we transform the longitude and latitude data as follows: $\rho\cdot\text{coord}+(1-\rho)\cdot\epsilon$ where $\rho=0.5$ and $\epsilon\sim\mathcal{N}(0,5)$.
@@
While the fit certainly is not perfect, the results do indicate that the random projection contains representations that are useful for the task at hand. Thus, this simple example illustrates that meaningful target representations should be recoverable from a sufficiently large latent space, given the projection of a small number of highly correlated features. Similarly, Alain & Bengio (2016) observe that even before training a convolutional neural network on MNIST data, the layer-wise activations can already be used to perform binary classification. In fact, it is well-known that random projections can be used for prediction tasks (Dasgupta, 2013).
This first experiment—and indeed the practice of probing LLMs that have seen vast amounts of data—can be seen as a form of inverse problem and common caveats such as non-uniqueness and instability apply (Haltmeier & Nguyen, 2023). Regularization can help mitigate these caveats (OM, 2001), but we confess that we did not carefully consider the parameter choice for $\lambda$, nor has this been carefully studied in the related literature to the best of our knowledge.
@@ 3.2 PCA as a Yield Curve Interpreter
We now move to a concrete application domain: Economics. Here, the yield curve, plotting the yields of bonds against their maturities, is a popular tool for investors and economists to gauge the health of the economy. The yield curve’s slope is often used as a predictor of future economic activity: a steep yield curve is associated with a growing economy, while a flat or inverted yield curve is associated with a contracting economy. To leverage this information in downstream modelling tasks, economists regularly use PCA to extract a low-dimensional projection of the yield curve that captures relevant variation in the data (e.g. Berardi & Plazzi (2022), Kumar (2022) and Crump & Gospodinov ).
@@ 3.3 LLMs for Economic Sentiment Prediction
Closely following the approach in Gurnee & Tegmark (2023b), we apply it to the novel Trillion Dollar Words (Shah et al., 2023) financial dataset, containing a curated selection of sentences formulated and communicated to the public by the Fed through speeches, meeting minutes and press conferences. (Shah et al., 2023) use this dataset to train a set of LLMs and rule-based models to classify sentences as either ‘dovish’, ‘hawkish’ or ‘neutral’. In the context of central banking, ‘hawkishness’ is typically associated with tight monetary policy: in other words, a ‘hawkish’ stance on policy favors high interest rates to limit the supply of money and thereby control inflation. The authors first manually annotate a sub-sample of the available data and then fine-tune various models for the classification task. Their model of choice, FOMC-RoBERTa (a fine-tuned version of RoBERTa (Liu et al., 2019)), achieves an $F_{1}$ score of around $>0.7$ on the test data. To illustrate the potential usefulness of the learned classifier, they use predicted labels for the entire dataset to compute an ad-hoc, count-based measure of ‘hawkishness’. This measure is shown to correlate with key economic indicators in the expected direction: when inflationary pressures rise, the measured level of ‘hawkishness’ increases, as central bankers react by raising interest rates to bring inflation back to target.
@@ 3.3.1 Linear Probes
For each layer, we compute linear probes through Ridge regression on two inflation indicators (the Consumer Price Index (CPI) and the Producer Price Index (PPI)) and US Treasury yields at different levels of maturity. To allow comparison with Shah et al. (2023), we let yields enter the regressions in levels. To measure price inflation we use percentage changes proxied by log differences. To mitigate issues related to over-parameterization, we follow the recommendation in Alain & Bengio (2016) to first reduce the dimensionality of the computed activations. In particular, we restrict our linear probes to the first 128 principal components of the embeddings of each layer. To account for stochasticity, we use an expanding window scheme with 5 folds for each indicator and layer. To avoid look-ahead bias, PCA is always computed on the sub-samples used for training the probe.
Figure 3 shows the out-of-sample root mean squared error (RMSE) for the linear probe, plotted against FOMC-RoBERTa’s $n$-th layer. The values correspond to averages across cross-validation folds. Consistent with related work (Alain & Bengio, 2016; Gurnee & Tegmark, 2023b), we observe that model performance tends to be higher for layers near the end of the transformer model. Curiously, for yields at longer maturities, we find that performance eventually deteriorates for the very final layers. We do not observe this for the training data, so we attribute this to overfitting.
@@
Figure 3: Out-of-sample root mean squared error (RMSE) for the linear probe plotted against FOMC-RoBERTa’s $n$-th layer for different indicators. The values correspond to averages computed across cross-validation folds, where we have used an expanding window approach to split the time series. As expected, model performance tends to be higher (average prediction errors are lower) for layers near the end of the transformer model.
@@ 3.3.2 Sparks of economic understanding?
Even though FOMC-RoBERTa, which is substantially smaller than the models tested in Gurnee & Tegmark (2023b), was not explicitly trained to uncover associations between central bank communications and the level of consumer prices, it appears that the model has distilled representations that can be used to predict inflation (although they certainly will not win any forecasting competitions). So, have we uncovered further evidence that LLMs “aren’t mere stochastic parrots”? Has FOMC-RoBERTa developed an intrinsic ‘understanding’ of the economy just by ‘reading’ central bank communications? Thus, can economists readily adopt FOMC-RoBERTa as a domain-relevant tool?
@@ 4 Human Proneness to Over-Interpretation
Linear probes and related tools from mechanistic interpretability were proposed in the context of monitoring models and diagnosing potential problems (Alain & Bengio, 2016). Favorable outcomes from probes merely indicate that the model “has learned information relevant for the property [of interest]” (Belinkov, 2021). Our examples demonstrate that this is achievable even for small models, while these have certainly not developed intrinsic “understanding” of the world. Thus, we argue that more conservative and rigorous tests for emerging capabilities of AI model are needed.
Generally, humans are prone to seek patterns everywhere. Meaningful patterns have proven useful in helping us make sense of our past, navigate our present and predict the future. Although this tendency to perceive patterns likely leads to evolutionary benefits even when the perceived patterns are false (Foster & Kokko, 2009), psychology has revealed a host of situations in which the ability to perceive patterns severely misfires, leading to irrational beliefs in the power of superstitions (Foster & Kokko, 2009), conspiracy theories (Van Prooijen et al., 2018), the paranormal (Müller & Hartmann, 2023), gambler’s fallacies (Ladouceur et al., 1996) and ‘pseudo-profound bullshit’ (Walker et al., 2019).